### v2

@@ Abstract
Zero-shot anomaly detection (ZSAD) enables identifying and localizing defects in unseen categories by relying solely on generalizable features rather than requiring any labeled examples of anomalies. However, existing ZSAD methods, whether using fixed or learned prompts, struggle under domain shifts because their training data are derived from limited training domains and fail to generalize to new distributions. In this paper, we introduce PILOT, a framework designed to overcome these challenges through two key innovations: (1) a novel dual-branch prompt learning mechanism that dynamically integrates a pool of learnable prompts with structured semantic attributes, enabling the model to adaptively weight the most relevant anomaly cues for each input image; and (2) a label-free test-time adaptation strategy that updates the learnable prompt parameters using high-confidence pseudo-labels from unlabeled test data. Extensive experiments on 13 industrial and medical benchmarks demonstrate that PILOT achieves state-of-the-art performance in both anomaly detection and localization under domain shift.
Accepted at BMVC 2025.
@@ 1 Introduction
Anomaly detection and localization is critical in safety-critical applications, including industrial inspection and medical diagnostics, where timely identification of rare deviations from normal patterns can prevent costly failures and improve patient outcomes. Traditional anomaly detection methods include supervised approaches that require labeled anomalous examples and unsupervised techniques relying solely on normal data [Bozic et al.(2021)Bozic, Tabernik, and Skocaj, Hojjati and Armanfard(2023), Li et al.(2021)Li, Sohn, Yoon, and Pfister, Roth et al.(2022)Roth, Pemula, Zepeda, Schölkopf, Brox, and Gehler, Karami et al.(2025)Karami, Ho, and Armanfard, Lai et al.(2023)Lai, Ho, and Armanfard]. Supervised methods require scarce and diverse anomaly labels, whereas unsupervised approaches avoid any need for anomalous samples by modeling only normal data, though they depend on comprehensive, representative normal data from the target domain [Chandola et al.(2009)Chandola, Banerjee, and Kumar]. Zero-shot Anomaly Detection (ZSAD) has emerged as a powerful approach by leveraging pretrained vision–language models (VLMs) such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], which learns joint representations of images and text from large-scale image–text data, to identify anomalies in visual data without requiring any target-domain annotations of normal or anomalous samples.
@@
Early successful ZSAD approaches utilized CLIP by designing explicit textual prompts that described both the object category and its condition, such as “normal bottle” or “anomalous transistor”. By crafting a large number of prompts that described relevant combinations of object category and condition, these methods enabled zero-shot classification and significantly outperformed vanilla CLIP baselines [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer, Chen et al.(2024)Chen, Zhang, Tian, He, Zhang, Wang, Wang, and Liu, Esmaeilpour et al.(2022)Esmaeilpour, Liu, Robertson, and Shu]. These approaches are known as fixed prompt methods because they rely on manually designed textual prompts. More recent methods have shifted towards a single learnable prompt, where the prompt itself is represented as a trainable embedding and optimized during training. This optimization is performed using auxiliary data, which refers to anomaly detection datasets that are separate from the target domain. For example, AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen] learns single object agnostic prompt embeddings to capture generic anomaly patterns, while AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi] further integrates visual knowledge into the prompt embeddings to enhance model adaptation.
@@
Despite these advances, both fixed and single learnable prompt approaches remain vulnerable to domain shift, which refers to performance degradation when data distribution in the target domain differs significantly from training data. Although pretrained VLMs like CLIP have strong generalization ability, their performance can drop sharply when there are significant differences between the data seen during pretraining and the target environment [Bose et al.(2024)Bose, Jha, Fini, Singha, Ricci, and Banerjee, Feng et al.(2024)Feng, Yu, Jin, Yu, Xiao, and Chen]. Even when learnable prompt embeddings are further trained on auxiliary anomaly detection datasets, that are different from the actual deployment domain, these models often struggle to represent the full diversity of possible anomalies or adapt to new categories encountered at test time. This limited generalization occurs for two main reasons: first, the auxiliary and target datasets may have significant distribution differences, as we empirically demonstrate in Appendix B.2; second, single prompt embeddings risk overfitting to the patterns and features specific to the auxiliary dataset, which further reduces their ability to handle unseen anomalies. As a result, prompts derived from narrowly defined auxiliary sources may fail to capture diverse visual features in the target data [Marzullo and Ranzini(2024)]. Therefore, to address these challenges, we introduce a framework that learns and dynamically weights multiple prompts, rather than relying on a fixed or single learnable prompt.
@@
On the other hand, Test-Time Adaptation (TTA) techniques have demonstrated improved model robustness in various computer vision tasks by refining model parameters on unlabeled test data [Nado et al.(2020)Nado, Padhy, Sculley, D’Amour, Lakshminarayanan, and Snoek, Liang et al.(2020)Liang, Hu, and Feng, Schneider et al.(2020)Schneider, Rusak, Eck, Bringmann, Brendel, and Bethge]. Although TTA is designed to address domain shift between training and test data, applying naive TTA strategies – without tailoring them to the specific task at hand – can inadvertently degrade performance. When used blindly, especially in tasks such as detection or localization, the adaptation process may optimize surrogate objectives that are misaligned with the true goals, often leading to suboptimal outcomes [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Liang et al.(2025)Liang, He, and Tan]. AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] was among the first to introduce TTA for anomaly localization refinement in ZSAD scenarios, but its focus is limited to localization and relies on fixed prompts. These limitations motivate us to explore how both anomaly detection and localization can be improved under domain shift by effectively utilizing unlabeled target data.  [^1] We present the Related Works section in Section A of the Supplementary Material..
@@
To address these challenges, we propose a novel framework, Prompt Learning with Integrated Label-Free Test-Time Adaptation for ZerO-ShoT Anomaly Detection (PILOT), a framework designed to mitigate the limitations of current ZSAD methods under domain shift. Instead of relying on a single learnable prompt, PILOT constructs a pool of learnable prompts, allowing the prompts to adaptively learn from various anomalies present in the auxiliary dataset. This design enables the model to adaptively weight the most relevant prompts for each input image, allowing for more nuanced representations of diverse anomalies and supporting more effective TTA. During inference, our label-free test-time adaptation strategy updates all learnable prompts according to their alignment with each test image, using high-confidence pseudo-labels generated from unlabeled target data. This targeted adaptation facilitates rapid and reliable adjustment to new domains (i.e., the test domain) without requiring any labeled data. In addition, to mitigate overfitting to the auxiliary dataset and enhance adaptation stability, PILOT incorporates an attribute memory bank, which consists of a curated set of fixed prompts that describe normal and anomalous states, leveraging pretrained semantic knowledge from the text encoder. This attribute memory bank serves as a semantic anchor, helping to stabilize localization performance during TTA. Together, these innovations enable PILOT to achieve robust anomaly detection and localization across a variety of challenging real-world scenarios. Our contributions are summarized as follows:
@@
  • To our knowledge, this is the first label-free TTA strategy in ZSAD that leverages high-confidence pseudo-labels from unlabeled test samples to rapidly adapt prompt parameters at inference, yielding improvements in both anomaly detection and localization under domain shift.
@@ 2.1 Learnable Prompt Pool
A successful recent approach for CLIP-based ZSAD employs a single learnable prompt, defined as a pair of sequences $(p^{n},p^{a})$ representing normal and anomalous states, respectively, for binary anomaly classification [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen]. Formally, this prompt can be represented as:
@@
where $L$ is the number of learnable prefix vectors, and each $v^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector trained jointly with the model. These vectors form a “soft prompt,” intuitively acting as special tokens whose values steer the model’s attention without modifying its core parameters, follows the prefix-tuning paradigm, which guides VLMs while keeping the main model weights frozen [Li and Liang(2021), Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan]. The placeholders “[normal state]” and “[anomalous state]” filled with descriptors of the object’s condition, such as “flawless” and “damaged.”
@@
Although the text encoder $f_{\mathrm{text}}$ is frozen to preserve CLIP’s rich pretrained textual representations, we employ a prefix-tuning strategy to enable effective fine-tuning and adaptation of CLIP. Inspired by similar prefix-tuning methods [Khattak et al.(2023)Khattak, Rasheed, Maaz, Khan, and Khan], we introduce additional learnable parameters to $f_{\mathrm{text}}$. More specifically, as is shown in Figure 1, layer-wise prefix vectors are defined as $\bigl{\{}\,w_{j}^{(\ell)}\in\mathbb{R}^{d}\mid j=1,\dots,J,\;\ell=1,\dots,L^{\prime}\bigr{\}}$, where $f_{\mathrm{text}}$ consists of a total of $J^{\prime}$ transformer layers, and $J\leq J^{\prime}$ denotes the number of initial layers to which learnable vectors are appended. $L^{\prime}$ represents the number of learnable vectors per layer. Specifically, these prefix vectors are prepended as additional tokens to the input sequence at the first $J$ transformer layers, leaving the original transformer layers unchanged. At each layer $j$, these vectors $\{w_{j}^{(\ell)}\}_{\ell=1}^{L^{\prime}}$ (illustrated as green blocks within the $f_{\mathrm{text}}$ module of Figure 1) are added before the standard, frozen token embeddings (represented as gray blocks in the the $f_{\mathrm{text}}$ module of Figure 1) prior to the self-attention mechanism in in each frozen transformer layer. This approach enables adaptation of the model’s internal representations without modifying the pretrained CLIP parameters, thereby preserving their robustness and generalization capabilities. Furthermore, as is illustrated in Figure 1, PILOT fine-tunes only the final $d$-dimensional projection layer within the vision encoder $f_{\mathrm{visual}}$. By restricting updates exclusively to this projection layer, we substantially reduce memory overhead and computational complexity. This targeted fine-tuning strategy also mitigates the risk of overfitting to auxiliary datasets, maintaining the strong generalization properties inherent to CLIP [Fahes et al.(2024)Fahes, Vu, Bursuc, Pérez, and de Charette].
@@ 2.2 Attribute Memory Bank
Although the learnable prompt pool $\mathcal{P}$ aims to provide adaptive and diverse representations tailored to the auxiliary dataset, it may still be susceptible to overfitting when deployed on target domains. To mitigate this issue, we further introduce a fixed attribute memory bank comprising attribute embeddings extracted from the pretrained CLIP text encoder $f_{\text{text}}$ (prior to applying our modifications). Recent works, such as WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer], have demonstrated the robustness of fixed textual prompts by encoding multiple natural-language templates. Inspired by this, we construct our attribute memory bank from a set of descriptive natural-language templates (e.g., “A bright photo of a [State] object”), where [State] indicates descriptive attributes of the object’s condition (e.g., “normal", “corroded", or “contaminated").
@@
where $\odot$” denotes element-wise multiplication, and $\mathrm{sim}(\cdot,\cdot)$ is the cosine similarity. Here, $u_{c}$ itself serves as the reference embedding for each attribute, as it is obtained directly from the pretrained $f_{\mathrm{text}}$ and already contains rich semantic information. Each score $\beta_{c}$ quantifies the alignment between the image and the $c$-th attribute embedding. The resulting scores $\{\beta_{c}\}_{c=1}^{C}$ quantify the alignment between the input and each attribute. Following [Jang et al.(2016)Jang, Gu, and Poole], we inject i.i.d. Gumbel noise into the attribute logits and apply the $\operatorname{entmax}{1.5}$ operator [Peters et al.(2019)Peters, Niculae, and Martins] to obtain a sparse yet differentiable weighting over the most prominent cues:
@@ 2.3 Prompt Aggregation via Orthogonal Projection
Following [Yang et al.(2021)Yang, He, Fan, Shi, Xue, Li, Ding, and Huang], to robustly leverage both the stable semantics from the attribute embedding $T_{att}$ and the adaptive semantics from the learnable prompts ($T_{n}$ for normal and $T_{a}$ for anomalous), we aggregate these representations using an orthogonal projection operation instead of simple averaging or concatenation. For each prompt embedding $T_{(\cdot)}$ (where $(\cdot)\in\{n,a\}$), we decompose it into components parallel and orthogonal to $T_{att}$:
@@ 2.4 Training Objectives
where $\sigma(\cdot)$ denotes the sigmoid function, $\mathrm{sim}(\cdot,\cdot)$ is cosine similarity, and $\mathrm{BCE}$ is the binary cross-entropy loss comparing the predicted score $\hat{y}$ to the ground-truth image label $y\in\{0,1\}$. For pixel-level localization, following [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen], we extract patch features $\{\,X^{(m)}\in\mathbb{R}^{H\times W\times d}\}_{m\in\mathcal{M}}$ from a set $\mathcal{M}$ of selected layers in $f_{\mathrm{visual}}$. Each $X^{(m)}$ is formed by reshaping the patch token sequence into a spatial grid of size $H\times W$ with $d$-dimensional embeddings. For each $m\in\mathcal{M}$, we compute an anomaly map using the aggregated text embedding:
@@
where $\hat{G}^{(m)}\in[0,1]^{H\times W}$ is the predicted anomaly probability map for layer $s$, $G$ is the ground-truth pixel-level mask, $\mathrm{Focal}(\cdot,\cdot)$ denotes the focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár], and $\mathrm{Dice}(\cdot,\cdot)$ denotes the Dice loss [Sudre et al.(2017)Sudre, Li, Vercauteren, Ourselin, and Jorge Cardoso]. This combination balances pixel-wise discrimination and overlap accuracy. The final training objective combines both terms $\mathcal{L}_{\mathrm{PILOT}}=\mathcal{L}_{\mathrm{image}}+\mathcal{L}_{\mathrm{pixel}}.$

### v3

@@ Abstract
Zero‑shot anomaly detection (ZSAD) enables identifying and localizing defects in unseen categories by relying solely on generalizable features rather than requiring any labeled examples of anomalies. However, existing ZSAD methods, whether using fixed or learned prompts, struggle under domain shifts because their training data are derived from limited training domains and fail to generalize to new distributions. In this paper, we introduce PILOT, a framework designed to overcome these challenges through two key innovations: (1) a novel dual-branch prompt learning mechanism that dynamically integrates a pool of learnable prompts with structured semantic attributes, enabling the model to adaptively weight the most relevant anomaly cues for each input image; and (2) a label-free test-time adaptation strategy that updates the learnable prompt parameters using high-confidence pseudo-labels from unlabeled test data. Extensive experiments on 13 industrial and medical benchmarks demonstrate that PILOT achieves state-of-the-art performance in both anomaly detection and localization under domain shift.
@@ 1 Introduction
Anomaly detection and localization is critical in safety-critical applications, including industrial inspection and medical diagnostics, where timely identification of rare deviations from normal patterns can prevent costly failures and improve patient outcomes. Traditional anomaly detection methods include supervised approaches that require labeled anomalous examples and unsupervised techniques relying solely on normal data [bovzivc2021mixed, hojjati2023dasvdd, li2021cutpaste, roth2022towards, karami2025graph, lai2023open]. Supervised methods require scarce and diverse anomaly labels, whereas unsupervised approaches avoid any need for anomalous samples by modeling only normal data, though they depend on comprehensive, representative normal data from the target domain [chandola2009anomaly]. Zero‑shot Anomaly Detection (ZSAD) has emerged as a powerful approach by leveraging pretrained vision–language models (VLMs) such as CLIP [clip], which learns joint representations of images and text from large-scale image–text data, to identify anomalies in visual data without requiring any target-domain annotations of normal or anomalous samples.
@@
Early successful ZSAD approaches utilized CLIP by designing explicit textual prompts that described both the object category and its condition, such as “normal bottle” or “anomalous transistor”. By crafting a large number of prompts that described relevant combinations of object category and condition, these methods enabled zero-shot classification and significantly outperformed vanilla CLIP baselines [winclip, clipad, zoc]. These approaches are known as fixed prompt methods because they rely on manually designed textual prompts. More recent methods have shifted towards a single learnable prompt, where the prompt itself is represented as a trainable embedding and optimized during training. This optimization is performed using auxiliary data, which refers to anomaly detection datasets that are separate from the target domain. For example, AnomalyCLIP [AnomalyCLIP] learns single object agnostic prompt embeddings to capture generic anomaly patterns, while AdaCLIP [AdaCLIP] further integrates visual knowledge into the prompt embeddings to enhance model adaptation.
@@
Despite these advances, both fixed and single learnable prompt approaches remain vulnerable to domain shift, which refers to performance degradation when data distribution in the target domain differs significantly from training data. Although pretrained VLMs like CLIP have strong generalization ability, their performance can drop sharply when there are significant differences between the data seen during pretraining and the target environment [bose2024stylip, feng2024rethinking]. Even when learnable prompt embeddings are further trained on auxiliary anomaly detection datasets, that are different from the actual deployment domain, these models often struggle to represent the full diversity of possible anomalies or adapt to new categories encountered at test time. This limited generalization occurs for two main reasons: first, the auxiliary and target datasets may have significant distribution differences, as we empirically demonstrate in Appendix B.2; second, single prompt embeddings risk overfitting to the patterns and features specific to the auxiliary dataset, which further reduces their ability to handle unseen anomalies. As a result, prompts derived from narrowly defined auxiliary sources may fail to capture diverse visual features in the target data [marzullo2024exploring]. Therefore, to address these challenges, we introduce a framework that learns and dynamically weights multiple prompts, rather than relying on a fixed or single learnable prompt.
@@
On the other hand, Test-Time Adaptation (TTA) techniques have demonstrated improved model robustness in various computer vision tasks by refining model parameters on unlabeled test data [nado2020evaluating, liang2020we, schneider2020improving]. Although TTA is designed to address domain shift between training and test data, applying naive TTA strategies – without tailoring them to the specific task at hand – can inadvertently degrade performance. When used blindly, especially in tasks such as detection or localization, the adaptation process may optimize surrogate objectives that are misaligned with the true goals, often leading to suboptimal outcomes [sun2020test, liu2021ttt, liang2025comprehensive]. AnovL [deng2023anovl] was among the first to introduce TTA for anomaly localization refinement in ZSAD scenarios, but its focus is limited to localization and relies on fixed prompts. These limitations motivate us to explore how both anomaly detection and localization can be improved under domain shift by effectively utilizing unlabeled target data.  [^1] We present the Related Works section in Section A of the Supplementary Material..
@@
To address these challenges, we propose a novel framework, Prompt Learning with Integrated Label‑Free Test‑Time Adaptation for ZerO‑ShoT Anomaly Detection (PILOT), a framework designed to mitigate the limitations of current ZSAD methods under domain shift. Instead of relying on a single learnable prompt, PILOT constructs a pool of learnable prompts, allowing the prompts to adaptively learn from various anomalies present in the auxiliary dataset. This design enables the model to adaptively weight the most relevant prompts for each input image, allowing for more nuanced representations of diverse anomalies and supporting more effective TTA. During inference, our label-free test-time adaptation strategy updates all learnable prompts according to their alignment with each test image, using high-confidence pseudo-labels generated from unlabeled target data. This targeted adaptation facilitates rapid and reliable adjustment to new domains (i.e., the test domain) without requiring any labeled data. In addition, to mitigate overfitting to the auxiliary dataset and enhance adaptation stability, PILOT incorporates an attribute memory bank, which consists of a curated set of fixed prompts that describe normal and anomalous states, leveraging pretrained semantic knowledge from the text encoder. This attribute memory bank serves as a semantic anchor, helping to stabilize localization performance during TTA. Together, these innovations enable PILOT to achieve robust anomaly detection and localization across a variety of challenging real-world scenarios. Our contributions are summarized as follows:
@@
  • To our knowledge, this is the first label‑free TTA strategy in ZSAD that leverages high‑confidence pseudo‑labels from unlabeled test samples to rapidly adapt prompt parameters at inference, yielding improvements in both anomaly detection and localization under domain shift.
@@ 2.1 Learnable Prompt Pool
A successful recent approach for CLIP-based ZSAD employs a single learnable prompt, defined as a pair of sequences $(p^{n},p^{a})$ representing normal and anomalous states, respectively, for binary anomaly classification [AnomalyCLIP]. Formally, this prompt can be represented as:
@@
where $L$ is the number of learnable prefix vectors, and each $v^{(i)}\in\mathbb{R}^{d}$ is a $d$-dimensional vector trained jointly with the model. These vectors form a “soft prompt,” intuitively acting as special tokens whose values steer the model’s attention without modifying its core parameters, follows the prefix-tuning paradigm, which guides VLMs while keeping the main model weights frozen [li2021prefix, khattak2023maple]. The placeholders “[normal state]” and “[anomalous state]” filled with descriptors of the object’s condition, such as “flawless” and “damaged.”
@@
Although the text encoder $f_{\mathrm{text}}$ is frozen to preserve CLIP’s rich pretrained textual representations, we employ a prefix-tuning strategy to enable effective fine-tuning and adaptation of CLIP. Inspired by similar prefix-tuning methods [khattak2023maple], we introduce additional learnable parameters to $f_{\mathrm{text}}$. More specifically, as is shown in Figure 1, layer-wise prefix vectors are defined as $\bigl\{\,w_{j}^{(\ell)}\in\mathbb{R}^{d}\mid j=1,\dots,J,\;\ell=1,\dots,L^{\prime}\bigr\}$, where $f_{\mathrm{text}}$ consists of a total of $J^{\prime}$ transformer layers, and $J\leq J^{\prime}$ denotes the number of initial layers to which learnable vectors are appended. $L^{\prime}$ represents the number of learnable vectors per layer. Specifically, these prefix vectors are prepended as additional tokens to the input sequence at the first $J$ transformer layers, leaving the original transformer layers unchanged. At each layer $j$, these vectors $\{w_{j}^{(\ell)}\}_{\ell=1}^{L^{\prime}}$ (illustrated as green blocks within the $f_{\mathrm{text}}$ module of Figure 1) are added before the standard, frozen token embeddings (represented as gray blocks in the the $f_{\mathrm{text}}$ module of Figure 1) prior to the self-attention mechanism in in each frozen transformer layer. This approach enables adaptation of the model’s internal representations without modifying the pretrained CLIP parameters, thereby preserving their robustness and generalization capabilities. Furthermore, as is illustrated in Figure 1, PILOT fine-tunes only the final $d$-dimensional projection layer within the vision encoder $f_{\mathrm{visual}}$. By restricting updates exclusively to this projection layer, we substantially reduce memory overhead and computational complexity. This targeted fine-tuning strategy also mitigates the risk of overfitting to auxiliary datasets, maintaining the strong generalization properties inherent to CLIP [fahes2024fine].
@@ 2.2 Attribute Memory Bank
Although the learnable prompt pool $\mathcal{P}$ aims to provide adaptive and diverse representations tailored to the auxiliary dataset, it may still be susceptible to overfitting when deployed on target domains. To mitigate this issue, we further introduce a fixed attribute memory bank comprising attribute embeddings extracted from the pretrained CLIP text encoder $f_{\text{text}}$ (prior to applying our modifications). Recent works, such as WinCLIP [winclip], have demonstrated the robustness of fixed textual prompts by encoding multiple natural-language templates. Inspired by this, we construct our attribute memory bank from a set of descriptive natural-language templates (e.g., “A bright photo of a [State] object”), where [State] indicates descriptive attributes of the object’s condition (e.g., “normal", “corroded", or “contaminated").
@@
where $\odot$” denotes element-wise multiplication, and $\mathrm{sim}(\cdot,\cdot)$ is the cosine similarity. Here, $u_{c}$ itself serves as the reference embedding for each attribute, as it is obtained directly from the pretrained $f_{\mathrm{text}}$ and already contains rich semantic information. Each score $\beta_{c}$ quantifies the alignment between the image and the $c$-th attribute embedding. The resulting scores $\{\beta_{c}\}_{c=1}^{C}$ quantify the alignment between the input and each attribute. Following [gumbel], we inject i.i.d. Gumbel noise into the attribute logits and apply the $\operatorname{entmax}{1.5}$ operator [entmax] to obtain a sparse yet differentiable weighting over the most prominent cues:
@@ 2.3 Prompt Aggregation via Orthogonal Projection
Following [yang2021dolg], to robustly leverage both the stable semantics from the attribute embedding $T_{att}$ and the adaptive semantics from the learnable prompts ($T_{n}$ for normal and $T_{a}$ for anomalous), we aggregate these representations using an orthogonal projection operation instead of simple averaging or concatenation. For each prompt embedding $T_{(\cdot)}$ (where $(\cdot)\in\{n,a\}$), we decompose it into components parallel and orthogonal to $T_{att}$:
@@ 2.4 Training Objectives
where $\sigma(\cdot)$ denotes the sigmoid function, $\mathrm{sim}(\cdot,\cdot)$ is cosine similarity, and $\mathrm{BCE}$ is the binary cross-entropy loss comparing the predicted score $\hat{y}$ to the ground-truth image label $y\in\{0,1\}$. For pixel-level localization, following [AnomalyCLIP], we extract patch features $\{\,X^{(m)}\in\mathbb{R}^{H\times W\times d}\}_{m\in\mathcal{M}}$ from a set $\mathcal{M}$ of selected layers in $f_{\mathrm{visual}}$. Each $X^{(m)}$ is formed by reshaping the patch token sequence into a spatial grid of size $H\times W$ with $d$-dimensional embeddings. For each $m\in\mathcal{M}$, we compute an anomaly map using the aggregated text embedding:
@@
where $\hat{G}^{(m)}\in[0,1]^{H\times W}$ is the predicted anomaly probability map for layer $s$, $G$ is the ground-truth pixel-level mask, $\mathrm{Focal}(\cdot,\cdot)$ denotes the focal loss [focal], and $\mathrm{Dice}(\cdot,\cdot)$ denotes the Dice loss [dice]. This combination balances pixel-wise discrimination and overlap accuracy. The final training objective combines both terms $\mathcal{L}_{\mathrm{PILOT}}=\mathcal{L}_{\mathrm{image}}+\mathcal{L}_{\mathrm{pixel}}.$